### v3

@@ 5.1 Implementation
For comparisons, we also implement the regularization and penalty-based methods. Thereby, an adaptation of Reg($\tau$), $\ell_{1}$ and $\ell_{\infty}$ Kim2020 ; Ralph2004 ; Nurkanovic2024b to the form of  (38) are solved in a homotopy loop with a decreasing parameter $\tau$. We initialize $\tau^{0}=1$ and update it using the rule $\tau^{k}=0.1\tau^{k-1}$. The loop terminates if either: the NLP solution is feasible and the complementarity tolerance $h_{\perp}(x)$ is below $10^{-9}$; NLP is locally infeasible; or a maximum number of $15$ iterations is reached. A homotopy loop is declared successful if it finds a stationary point.
Additionally, we compare MPECopt to solving the MPEC directly as an NLP in 2 with IPOPT, and solving its MINLP reformulation (21) with a nonlinear branch-and-bound method implemented in Bonmin Bonami2008 .
@@ 5.2 Results on MacMPEC benchmark
We compare MPECopt to other mentioned approaches. For Phase I of MPECopt, we use Algorithm (2) with Gurobi as LPEC solvers, as it had the highest success compared to other variants we tested. In the experiments, we compare in total six algorithms:
@@
We also tried the $\ell_{1}$ penalty method, which similar computation time as $\ell_{\infty}$, but a slightly lower success rate. Moreover using HiGHS instead of Gurobi lead to slightly slower method and marginally lower success rate.
Therefore, for better readability, we omit these variants from the plots. Since all method use IPOPT as NLP solver, we can directly compare the computation times. Fig. 5 reports the absolute and relative performance profiles for this experiment.
Note that, if MPECopt is successful, it always returns a B-stationary point, whereas for the other methods, we can only know that it is a B-stationary point if it is also S-stationary. To verify B-stationarity of the solutions computed by all other methods, we proceed as described in Appendix B.
We do not add the computational cost of verifying B-stationarity to the overall computation time of the other methods.
@@
The MPECopt method inherits the efficacy of the standard methods used in Phase I but can often be even faster, as it requires solving fewer NLPs. This is because, already for moderate values of $\tau$, the LPEC can identify a feasible BNLP in Phase I. Consequently, MPECopt needs to solve fewer NLPs than Reg and $\ell_{\infty}$. The branch-and-bound MINLP method solves the largest number of NLPs, which also explains its slow computation times.
@@
We prove in Section 4 that, unless we need to verify B-stationarity, the LPECs do not need to be solved to global optimality; finding a feasible point is sufficient. Moreover, Proposition 3.1 shows that if the solution is S-stationary, it can be verified by solving a single LP. We examine these theoretical results by plotting histograms of the number of nodes in a MILP call (corresponding to the number of LPs solved) in both Phase I and Phase II. Fig. 6a shows the histogram when the MILP is solved to global optimality, while Fig. 6b shows the case when it is terminated early, at a feasible point. In both cases, Phase II only needed to solve LPs, and in Phase I we observe a significant reduction in the number of LPs to be solved in several cases. In fact, with early termination, on MacMPEC we only need to solve LPs. This illustrates our theoretical results: early termination of the MILP method reduces the number of LPs that must be solved. This did not significantly affect the overall computation time, as the largest computational burden on the MacMPEC problem set came from solving the NLP subproblems.
@@ 5.3 Synthetic nonlinear MPEC benchmark
Similar to the results on MacMPEC (Fig. 5), we see that the MPECopt variants are the fastest and most successful algorithms even on the large problem instances. For the same reasons as in the previous experiments, they require fewer total NLP solves. Moreover, on this set, the differences in computation times between MPECOpt-Reg-Gurobi (96.16% success rate), MPECOpt-$\ell_{\infty}$-Gurobi (94.67% success rate), and Pen-$\ell_{\infty}$ (94.67% success rate) are more pronounced, whereas Reg (94.67% success rate) needs to solve more NLPs and is therefore slower. The lower success rates of the regularization- and penalty-based methods compared to their MPECopt counterparts are due to either reaching the maximum number of iterations or hitting the time limit. In other cases, the methods failed by converging to a point of local infeasibility, i.e., when the restoration phase failed. Directly solving the NLP is by far the least successful approach (16.00% success rate), and in most cases it timed out after reaching the 45-minute limit. This further justifies the development of tailored methods for solving MPECs. As shown in Fig. 7a, the other methods compute solutions considerably faster and rarely run into timeouts.
@@
In conclusion, we see that the proposed method is competitive with other existing approaches. Moreover, the experiments demonstrate that solving an LPEC was never a computational bottleneck, even for the larger problem instances. Significant savings can be achieved through early termination, in line with our theoretical predictions.
@@ 6 Conclusions and summary
This paper introduces MPECopt, a two-phase method for computing B-stationary points of mathematical programs with complementarity constraints (MPECs). The first phase identifies a feasible point of the MPEC or certifies local infeasibility. The second phase iterates over feasible points to find a B-stationary point by solving a finite sequence of branch nonlinear programs (BNLPs) and linear programs with complementarity constraints (LPECs). We provide a detailed theoretical analysis of the two LPEC subproblems and prove that the algorithm terminates in a finite number of iterations.
We also provide an open-source implementation of MPECopt, available at github.com/nosnoc/mpecopt. Our implementation leverages robust and mature NLP and MILP solvers. We tested MPECopt on two extensive numerical benchmarks: the MacMPEC set and a new synthetic nonlinear MPEC test set, containing problems with up to four thousand complementarity constraints. The numerical results confirm the efficacy of the proposed method for finding B-stationary points. MPECopt is generally faster and more reliable than existing methods, requiring fewer NLP solves than homotopy-based approaches.
@@ Acknowledgments
Appendix A Proofs of lemmata from Section 3
@@
$\begin{split}\Omega^{\mathrm{F}}=\Omega^{\mathrm{R}}&\bigcup_{i\in{\mathcal{I}}_{0+}^{k}}\{d\mid x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0\}\\ &\bigcup_{i\in{\mathcal{I}}_{+0}^{k}}\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}\\ &\bigcap_{i\in\bar{{\mathcal{A}}}^{k}}\{d\mid c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0\}.\end{split}$ (39)
@@
Now observe that, for example, for any $i\in{\mathcal{I}}_{+0}^{k}$ it holds that $x_{1,i}^{k}>0$ and $x_{2,i}=0$. If $\rho$ is chosen such that $d_{1,i}<\rho<x_{1,i}^{k}$ then $x_{1,i}^{k}+d_{1,i}>0$ and $0\leq x_{1,i}^{k}+d_{1,i}\perp d_{2,i}\geq 0$ can be only satisfied for $d_{2,{i}}=0$. Thus, the branch $\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}$ is infeasible. Similar reasoning holds for any $i\in{\mathcal{I}}_{0,+}^{k}$. Therefore, by picking $\rho<\bar{\rho}_{2}=\min\{\{x_{1,i}\mid i\in{\mathcal{I}}_{+0}^{k}\}\cup\{x_{2,i}\mid i\in{\mathcal{I}}_{0+}^{k}\}\}$, none of the terms in the unions in (39) is feasible. The equations (16) is follows from writing $\bar{\rho}=\min(\bar{\rho}_{1},\bar{\rho}_{2})$ more explicitly. Thus, for $\rho<\bar{\rho}$, we have $\Omega^{\mathrm{F}}\cap\{\|d\|_{\infty}\leq\rho\}=\Omega^{R}\cap\{\|d\|_{\infty}\leq\rho\}$. As the two LPECs have the same objective, their sets of local minimizers coincide. ∎
Proof of Lemma 3.2: It follows from Lemma 3.1 that for $\rho\in(0,\bar{\rho})$ both the full and reduced LPEC have the same feasible set.
Thus, we focus solely on the reduced LPEC. Recall from Eq. (10) that $\Omega=\bigcup_{(\mathcal{I}_{1}^{k},\mathcal{I}_{2}^{k})\in\mathcal{P}^{k}}\Omega_{\mathrm{BNLP}{(\mathcal{I}_{1}^{k},\mathcal{I}_{2}^{k})}}$, where $\mathcal{P}^{k}$ denotes the set of all possible partitions $(\mathcal{I}_{1}^{k},\mathcal{I}_{2}^{k})$ at $x^{k}\in\Omega$, cf. (7). This relation also holds for the tangent cones of the MPEC and the BNLPs, as well as for the MPEC-linearized feasible cone and the standard linearized feasible cones of the BNLPs (Flegel2005a, , Lemma 3.2). The latter reads explicitly as:
@@ Appendix C Comparison of different LPEC solvers
Somewhat surprisingly, the global MILP algorithms significantly outperform the local regularization and penalty-based methods. The MILP-based methods are up to two orders of magnitude faster than regularization-based methods for LPECs on the MacMPEC test set. This difference can be explained by several observations. The MILP methods solve the LPEC to global optimality (sufficient for verifying B-stationarity), whereas the others compute only a stationary point (which may not be a local minimum or B-stationary point). Consequently, LPECs solved by regularization-based methods are more often locally infeasible in Phase I or require solving more LPECs until a B-stationary point is verified in Phase II. This confirms the arguments in Section 3.2.2 on efficiently solving LPECs using MILP reformulations. Moreover, when the solution is S-stationary, solving the MILP requires solving a single LP, complying with Proposition 3.1.
This experiment suggests that solving LPECs with MILP methods as in MacMPEC can be done very efficiently in practice, both with commercial (Gurobi) and open-source (HiGHS) solvers.

### v4

@@ 5.1 Implementation
For comparisons, we also implement the regularization and penalty-based methods. Thereby, an adaptation of Reg($\tau$), $\ell_{1}$ and $\ell_{\infty}$ Kim2020 ; Ralph2004 ; Nurkanovic2024b to the form of  (38) are solved in a homotopy loop with a decreasing parameter $\tau$. We initialize $\tau^{0}=1$ and update it using the rule $\tau^{k}=0.1\tau^{k-1}$. The loop terminates if either: the NLP solution is feasible and the complementarity tolerance $h_{\perp}(x)$ is below $10^{-9}$; NLP is locally infeasible; or a maximum number of $15$ homotopy iterations is reached. A homotopy loop is declared successful if it finds a stationary point. Additionally, we compare MPECopt to solving the MPEC directly as an NLP in (2) with IPOPT, and solving its MINLP reformulation (21) with a nonlinear branch-and-bound method implemented in Bonmin Bonami2008 .
@@ 5.2 Results on MacMPEC benchmark
We compare MPECopt to other mentioned approaches. For Phase I of MPECopt, we use Algorithm 2 with Gurobi as LPEC solvers, as it had the highest success compared to other variants we tested. In the experiments, we compare in total six algorithms:
@@
We also tried the $\ell_{1}$ penalty method, which has a similar computation time as $\ell_{\infty}$, but a slightly lower success rate. Moreover using HiGHS instead of Gurobi lead to slightly slower method and marginally lower success rate. Therefore, for better readability, we omit these variants from the plots. Since all methods use IPOPT as the NLP solver, we can directly compare the computation times. Fig. 5 reports the absolute and relative performance profiles for this experiment.
Note that, if MPECopt is successful, it always returns a B-stationary point, whereas for the other methods, we can only know that it is a B-stationary point if it is also S-stationary. To verify B-stationarity of the solutions computed by all other methods, we proceed as described in Appendix B. We do not add the computational cost of verifying B-stationarity to the overall computation time of the other methods.
@@
The MPECopt method inherits the efficacy of the standard methods used in Phase I but can often be even faster, as it requires solving fewer NLPs. This is because, already for moderate values of $\tau$, the LPEC can identify a feasible BNLP in Phase I. Consequently, MPECopt needs to solve fewer NLPs than Reg and $\ell_{\infty}$ in most cases. The branch-and-bound MINLP method solves the largest number of NLPs, which also explains its slow computation times.
@@
We have shown in Section 4 that, unless we need to verify B-stationarity, the LPECs do not need to be solved to global optimality; finding a feasible point is sufficient. Moreover, Proposition 3.1 shows that if the solution is S-stationary, it can be verified by solving a single LP. We examine these theoretical results by plotting histograms of the number of nodes in a MILP call (corresponding to the number of LPs solved) in both Phase I and Phase II. Fig. 6a shows the histogram when the MILP is solved to global optimality, while Fig. 6b shows the case when it is terminated early, at a feasible point. In both cases, Phase II only needed to solve LPs, and in Phase I we observe a significant reduction in the number of LPs to be solved in several cases. In fact, with early termination, on MacMPEC we only need to solve LPs. This illustrates our theoretical results: early termination of the MILP method reduces the number of LPs that must be solved. This did not significantly affect the overall computation time, as the largest computational burden on the MacMPEC problem set came from solving the NLP subproblems.
@@ 5.3 Synthetic nonlinear MPEC benchmark
Similar to the results on MacMPEC (Fig. 5), we see that the MPECopt variants are the fastest and most successful algorithms even on the large problem instances. For the same reasons as in the previous experiments, they require fewer total NLP solves. Moreover, on this set, the differences in computation times between MPECOpt-Reg-Gurobi (96.16% success rate), MPECOpt-$\ell_{\infty}$-Gurobi (94.67% success rate), and Pen-$\ell_{\infty}$ (94.67% success rate) are more pronounced, whereas Reg (94.67% success rate) needs to solve more NLPs and is therefore slower. The lower success rates of the regularization- and penalty-based methods compared to their MPECopt counterparts are due to either reaching the maximum number of iterations or hitting the time limit. In other cases, the methods failed by converging to a point of local infeasibility, i.e., when the restoration phase failed. Directly solving the NLP is by far the least successful approach (16.00% success rate), and in most cases it timed out after reaching the 45-minute limit. This further justifies the development of tailored methods for solving MPECs. As shown in Fig. 7a, the MPEC-tailored methods compute solutions considerably faster and rarely run into timeouts.
@@
In conclusion, we can see that the MPECopt is competitive with other existing approaches. Unlike all other methods, it always provides a certificate of B-stationarity. Moreover, the experiments demonstrate that solving an LPEC was never a computational bottleneck, even for the larger problem instances. Significant savings can be achieved through early termination, in line with our theoretical predictions.
@@ 6 Conclusions and summary
This paper introduces MPECopt, a two-phase method for computing B-stationary points of mathematical programs with complementarity constraints (MPECs). The first phase identifies a feasible point of the MPEC or certifies local infeasibility. The second phase iterates over feasible points to find a B-stationary point by solving a finite sequence of branch nonlinear programs (BNLPs) and linear programs with complementarity constraints (LPECs). We perform a detailed convergence analysis of the algorithm.
We also provide an open-source implementation of MPECopt, available at github.com/nosnoc/mpecopt. The implementation leverages robust and mature NLP and MILP solvers. We tested MPECopt on two numerical benchmarks: the MacMPEC set and a new synthetic nonlinear MPEC test set containing problems with up to four thousand complementarity constraints. The numerical results confirm the efficacy of the proposed method for finding B-stationary points. MPECopt is generally faster and more reliable than existing methods, requiring fewer NLP solves than homotopy-based approaches.
@@ Acknowledgments
Data availability
All data necessary to reproduce the results in this paper are publicly available in the GitHub repository https://github.com/nosnoc/mpecopt. This repository contains the code, experiment scripts, and data used to generate the numerical results reported in the manuscript.
Competing interests
The authors declare no competing interests.
Appendix A Proofs of lemmas from Section 3
@@
$\begin{split}\Omega^{\mathrm{F}}=\Omega^{\mathrm{R}}&\bigcap_{i\in\bar{{\mathcal{A}}}^{k}}\{d\mid c_{i}(x^{k})+\nabla c_{i}(x^{k})^{\top}d\geq 0\}\\ &\bigcup_{i\in{\mathcal{I}}_{0+}^{k}}\{d\mid x_{1,i}^{k}+d_{1,i}\geq 0,\ x_{2,i}^{k}+d_{2,i}=0\}\\ &\bigcup_{i\in{\mathcal{I}}_{+0}^{k}}\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}.\end{split}$ (39)
@@
Now observe that, for example, for any $i\in{\mathcal{I}}_{+0}^{k}$ it holds that $x_{1,i}^{k}>0$ and $x_{2,i}=0$. If $\rho$ is chosen such that $d_{1,i}<\rho<x_{1,i}^{k}$ then $x_{1,i}^{k}+d_{1,i}>0$ and $0\leq x_{1,i}^{k}+d_{1,i}\perp d_{2,i}\geq 0$ can be only satisfied for $d_{2,{i}}=0$. Thus, the branch $\{d\mid x_{1,i}^{k}+d_{1,i}=0,\ x_{2,i}^{k}+d_{2,i}\geq 0\}$ cannot be reached, i.e., it is infeasible. Similar reasoning holds for any $i\in{\mathcal{I}}_{0,+}^{k}$. Therefore, by picking $\rho<\bar{\rho}_{2}=\min\{\{x_{1,i}\mid i\in{\mathcal{I}}_{+0}^{k}\}\cup\{x_{2,i}\mid i\in{\mathcal{I}}_{0+}^{k}\}\}$, none of the terms in the unions in (39) is feasible. The equation (16) follows from writing $\bar{\rho}=\min(\bar{\rho}_{1},\bar{\rho}_{2})$ explicitly. Thus, for $\rho<\bar{\rho}$, we have $\Omega^{\mathrm{F}}\cap\{\|d\|_{\infty}\leq\rho\}=\Omega^{R}\cap\{\|d\|_{\infty}\leq\rho\}$. As the two LPECs have the same objective, their sets of local minimizers coincide. ∎
Proof of Lemma 3.2: It follows from Lemma 3.1 that for $\rho\in(0,\bar{\rho})$ both the full and reduced LPEC have the same feasible set. Thus, we focus solely on the reduced LPEC. Denote by $\mathcal{P}^{k}$ the set of all possible partitions $(\mathcal{I}_{1}^{k},\mathcal{I}_{2}^{k})$ at $x^{k}\in\Omega$ (see Eq. (7) for definitions). The MPEC-MFCQ implies the MPEC-ACQ, hence the MPEC-linearized feasible cone and the standard linearized feasible cones of the BNLPs satisfy the following (Flegel2005a, , Lemma 3.1):
@@ Appendix C Comparison of different LPEC solvers
Somewhat surprisingly, the global MILP algorithms significantly outperform the local regularization and penalty-based methods. The MILP-based methods are up to two orders of magnitude faster than regularization-based methods for LPECs on the MacMPEC test set. This difference can be explained by several observations. The MILP methods solve the LPEC to global optimality (sufficient for verifying B-stationarity), whereas the others compute only a stationary point (which may not be a local minimum or B-stationary point). Consequently, LPECs solved by regularization-based methods are more often locally infeasible in Phase I or require solving more LPECs until a B-stationary point is verified in Phase II. This confirms the arguments in Section 3.2.2 on efficiently solving LPECs using MILP reformulations. Moreover, when the solution is S-stationary, solving the MILP requires solving a single LP, complying with Proposition 3.1. This experiment suggests that solving LPECs with MILP methods as in MacMPEC can be done very efficiently in practice, both with commercial (Gurobi) and open-source (HiGHS) solvers.